OpenStudy (anonymous):

sin theta= 3/4 located in quadrant 2 what is sin2theta, cos2theta, and tan2theta. Please help, have no clue how to even start

OpenStudy (anonymous):

\[\sin(2 \theta) = 2 \cos \theta \sin \theta\] \[\cos(2 \theta) = \cos^2 \theta - \sin^ 2 \theta \] \[\tan 2 \theta = \frac{ 2 \tan \theta }{ 1- \tan ^2 \theta }\] you can use these identities